OpenStudy (anonymous):

Least period of function? y=4sin t/5

OpenStudy (dumbcow):

for function \[y = A \sin (Bt)\] period = 2pi/B \[p = \frac{2\pi}{1/5} = 2\pi*5 = 10\pi\]